Set up a table of values (minimum 5 ordered pairs) equally arranged around the axis of symmetry x value to prepare to graph:
y=x^2-10x+25

Answer by stanbon(75887)

Set up a table of values (minimum 5 ordered pairs) equally arranged around the axis of symmetry x value to prepare to graph:
y=x^2-10x+25
-------------
(x-5)^2 = y-0
------------------------
Axis of symmetry: x = 5
--------------------------
Table:
If x = 3, y = 3^2-10*3+25 = 4
If x = 4, y = 4^2-10*4+25 = 1
If x = 6, y = 6^2-10*6+25 = 1
If x = 7, y = 7^2-10*7+25 = 4
==================================
Cheers,
Stan H.